OpenStudy (cagliero423):

URGENT: Can anyone help me with composition of functions? I'm stuck on a problem and can't move on until I figure it out.

OpenStudy (cagliero423):

Given f(x)=x^2+3 and g(x)= x+5/x find (g x f)(4)

OpenStudy (cagliero423):

I keep getting 171/76 and my answer choices are 15, -15, 14/19, 24/19

OpenStudy (cagliero423):

@Xspead can you help??

OpenStudy (anonymous):

No I don't totally understand this sorry :(

OpenStudy (cagliero423):

Oh, that's okay(x

OpenStudy (cagliero423):

Well I figured it out, it was 24/19 if anyone cares hahah